OpenStudy (anonymous):

lim x-> infinity ((sqrt(x^2-3x)-x) how do i find the limit? please help don't understand

OpenStudy (anonymous):

try multiplying by \[\frac{\sqrt{x^2-3x}+x}{\sqrt{x^2-3x}+x}\]

OpenStudy (anonymous):

i tried that gives me -3x/sqrt(x^2-3x)+x dont know what to do after

OpenStudy (anonymous):

that is -infinity/infinity+infinity

OpenStudy (anonymous):

but it needs to be a number right?

OpenStudy (anonymous):

you get \[\frac{-3x}{\sqrt{x^2-3x}+x}\] yes

OpenStudy (anonymous):

now you can pretty much do this by your eyeballs. ignore the \(-3x\) inside the radical in the denominator you get \[\frac{-3x}{\sqrt{x^2}+x}\] \[=\frac{-3x}{x+x}=\frac{-3x}{2x}=-\frac{3}{2}\]

OpenStudy (anonymous):

oh yeah thats right i understand tnx

OpenStudy (anonymous):

if that is not sufficient, you have to divide top and bottom by \(x\) be careful inside the radical

OpenStudy (anonymous):

yw